OpenStudy (cometailcane):

The density of gold in 19.3 g/cm3. Convert the density to kg/m3. Show your work. What volume of water would a 10.3 g piece of gold displace?

OpenStudy (aveline):

You'll want to use dimensional analysis for this problem. \[\frac{ 19.3gAu }{ cm^ {3} } \times \left( \frac{ 100cm }{ m }\right)^ {3} \times \frac{ kg }{ 1000g }\]

OpenStudy (aveline):

The density of the gold is 19.3g per centimeter cubed. We multiply by the conversion unit 100cm per m (there are 100 centimeters in a meter). Because we want to go to meters cubed from centimeters cubed, we cube this conversion unit. Then, we use kg/1000g (another conversion unit) to change the grams to kilograms.